# Exploring non-Euclidean photonics: Pseudosphere microlaser

### v1

@@ Appendix B Fourier transform versus Lomb-Scargle periodogram
A Discrete Fourier Transform (DFT) at the frequency $f$ can be written
@@
Therefore $k_{n+1}-k_{n}$ is not constant for the experimental spectra. In the top panel of Fig. 4(b), we plot the Fourier transform of the tractoid experimental spectrum of Fig. 4(a) calculated by the FFT algorithm implemented on Python in the scipy library, which corresponds to the modulus squared of formula (41) using the constant sampling $k_{n+1}-k_{n}=(k_{2}-k_{1})/N$. For comparison, in the middle panel of Fig. 4(b), we plot the modulus square of expression (41) using the actual $(k_{n})$, which are not equidistant. Note that the main peak is significantly shifted, revealing an artefact: the raw DFT (to panel) is peaked at a length which is longer than the perimeter (which is not consistent with WGM), while for in the middle panel the length is slightly shorter than the perimeter (in agreement with WGM).
The Lomb periodogram $P_{LS}(f)$ plotted in the bottom panel of Fig. 4(b) was computed by the lombscargle routine implemented on Python in the scipy library. Its expression is given in [19] and is similar to the formula in the seminal paper by [20]. It is peaked at the same length than the well-sampled DFT in the middle panel of Fig. 4(b). Moreover the signal-to-noise ratio is much better. Hence we prefer the Lomb periodogram for analyzing the experimental spectra.
@@ Appendix C Stability computations of periodic geodesics
This Appendix deals with the stability of geodesics on the tractoid surface. First we derive the monodromy matrix for propagation $\Pi$, then the monodromy matrix for reflection $\mathcal{R}$. Finally in App. C.3 and C.4, we apply these formulas to periodic geodesics of the full and truncated tractoid, respectively.
@@ C.1 Monodromy matrix for propagation
In this paragraph, we derive the monodromy matrix $\Pi(t)$ for the propagation along a geodesic on the tractoid surface. The usual $(\zeta,\sin(\chi))$ coordinate system is convenient where $\zeta$ is the coordinate in a direction orthogonal to the geodesic tangent vector and $\chi$ is the deviation from the original propagation direction, as shown in Fig. 15. The coordinate $\sin(\chi)$ corresponds to $d\zeta/dt$, ie. the local slope with respect to the unperturbed geodesic. The matrix $\Pi(t)$ is defined by the following system at the first order
$\begin{pmatrix}\zeta(t)\\ \zeta^{\prime}(t)\end{pmatrix}=\Pi(t)\begin{pmatrix}\zeta(0)\\ \zeta^{\prime}(0)\end{pmatrix}$ (42)
The Jabobi equation [11] describes the evolution of $\zeta$ along a given geodesic as a function of $t$ and can be written
@@
Deriving (44) and after identification with (42), one obtains
$\Pi(t)=\begin{pmatrix}\cosh\left(\frac{t}{R_{0}}\right)&R_{0}\sinh\left(\frac{t}{R_{0}}\right)\\ \frac{1}{R_{0}}\sinh\left(\frac{t}{R_{0}}\right)&\cosh\left(\frac{t}{R_{0}}\right)\end{pmatrix}$ (45)
which is the monodromy matrix for propagation along a geodesic on the tractoid. As expected, it has the property $\Pi(t_{1})\Pi(t_{2})=\Pi(t_{1}+t_{2})$ and it tends to the flat propagation matrix in the limit of vanishing Gaussian curvature $R_{0}\to\infty$. Note that $\det(\Pi)=1$ and $\operatorname{tr}(\Pi)=2\cosh(\frac{t}{R_{0}})$ which is larger than 2. Therefore, one of its eigenvalues is larger than 1 and the propagation is unstable as it is expected on a negatively curved surface.
Figure 15: Schematic for the computation of the propagation monodromy matrix. The black curve is the reference geodesic. It is parametrized by $t$ such that the black dot corresponds to $t=0$. The blue dotted line is the perturbated geodesic. The initial point at $t=0$ is represented by the blue dot and is shifted by a distance $\zeta(0)$. The initial angle of the blue curved with respect to the reference is called $\chi$
@@ C.2 Monodromy matrix for reflection
In this section, we derive the monodromy matrix for the reflection of a ray on the surface boundary. First, the general case of a given surface and a given boundary is studied. Then this result is applied to the specific case of the tractoid.
@@ C.2.1 The general case
The surface is defined by the vector $\vec{F}(u,v)$ [cf. Eqs. (1-3)] and parameterized by the variables $u$ and $v$. The boundary is a curve belonging to the surface and defined by the vector $\vec{c}\,(t)$. The parametrization is chosen such that $\|\frac{d\vec{c}}{dt}\|=1$. This choice simplifies the definition of the curvature of the curve which is then the norm of the second derivative: $\kappa=\|\frac{d^{2}\vec{c}}{dt^{2}}\|$.
@@
It remains to calculate the tangent curvature $\kappa_{t}$ because we must only consider the curvature in the tangent plane where the reflection takes place. To formalize that, we call $\vec{N}$ the unit vector normal to the surface. If the curvature of the boundary $\frac{d^{2}\vec{c}}{dt^{2}}$ is collinear to $\vec{N}$ then all the curvature is outside the tangent plane. Therefore, the tangent curvature $\kappa_{t}$ is zero. This is the case when the curve $\vec{c}$ is a geodesic. Hence, a reflection on a geodesic is equivalent to a reflection on a plane mirror in the 2D Euclidean case. In general, the vector normal to the surface $\vec{N}$ and the curvature of the boundary $\vec{c}$ are not collinear. The tangent curvature $\kappa_{t}$ is the norm of the projection of $\frac{d^{2}\vec{c}}{dt^{2}}$ onto the tangent plane:
@@ C.3 Stability computations for the full tractoid
In this paragraph, the full infinite tractoid is considered. It means that only the periodic geodesics described in Fig. 5 exist. Using the same notations as in Sec. IV, the total monodromy matrix for the $(M,P)$-periodic geodesic can be written
@@
with $\mathcal{M}_{0}(M,P)=\Pi(M,P)R(M,P)$ the product of the propagation matrix and the reflection matrix for the $(M,P)$-periodic geodesic. Combining formulas (13) and (14) with the expression of the matrices in Eqs. (15) and (16), one we obtain the following expressions for $\Pi(M,P)$
$\Pi(M,P)=\begin{pmatrix}\Pi_{1}&R_{0}\Pi_{2}\\ \frac{\Pi_{2}}{R_{0}}&\Pi_{1}\end{pmatrix}$
@@
and for $R(M,P)$
$R(M,P)=\begin{pmatrix}-1&0\\ \frac{2}{R_{0}}\frac{M}{P\pi}\left(1+\left(\frac{P\pi}{M}\right)^{2}\right)^{\frac{1}{2}}&-1\end{pmatrix}$ (56)
@@ C.4 Stability computations for the truncated tractoid
However, rays can also be reflected on the top boundary and therefore a new type of periodic geodesic appears. Since the top boundary is also a straight horizontal line, the reflection angles are going to be preserved. It entails that such periodic geodesics are composed of a series of identical patterns, made of arc circles. Each such periodic geodesic is uniquely identified by a pair of positive relatively prime integers that will also be called $(M,P)$, $M$ being the number of reflections on the lower boundary and $P$ the rotation number. The $(4,3)$ periodic geodesic is sketched in Fig. 16.
@@
As previously, the total monodromy matrix of the $(M,P)$ periodic geodesics $\mathcal{M}^{\prime}(M,P)$ can be written
@@
$\mathcal{M}^{\prime}_{0}(M,P)=\Pi(M,P)\mathcal{R}(M,P)\Pi(M,P)\tilde{\mathcal{R}}(M,P)$ (61)
where $\Pi(M,P)$ accounts for the propagation from the upper boundary to the lower one or the opposite, $\mathcal{R}(M,P)$ for the reflection at the boundary $y=1$ and ~$\tilde{\mathcal{R}}(M,P)$ for the reflection at the boundary $y=y_{max}$. Note that $\epsilon=+1$ in $\mathcal{R}(M,P)$, and $\epsilon=-1$ in ~$\tilde{\mathcal{R}}(M,P)$.
After cumbersome but trivial calculations, the trace of $M^{\prime}_{0}(M,P)$ is equal to 2. These periodic geodesics are thus also marginally stable as expected.
@@ Appendix E Matrices for solving the scalar wave equation
The following matrices are used in Sec. V.4 to solve the scalar wave equation via an eigenvalue problem:
$T=\frac{1}{h^{2}}\begin{pmatrix}2&-1&0&\dots&0\\ -1&2&-1&&\\ &&\ddots&&\\ &&-1&2&-1\\ 0&\dots&0&-1&2\end{pmatrix}$ (71)
$V=m^{2}\begin{pmatrix}e^{2\eta_{1}}&&\\ &\ddots&\\ &&e^{2\eta_{N}}\end{pmatrix}$ (72)
@@ Appendix F Weyl law
The integrated average density of modes is basically to the number of eigenmodes from 00 to $x:=kR_{0}$ and is given by the Wey law [4]:
$N_{Weyl}(x)=\frac{\mathcal{A}}{2\pi}x^{2}-\frac{\mathcal{P}}{4\pi}x-\frac{\mathcal{A}}{12\pi}$ (73)
where $\mathcal{P}$ and $\mathcal{A}$ are the perimeter of the tractoid in units of $R_{0}$, and the area of the tractoid in units of $R_{0}^{2}$, respectively. For the full tractoid (i.e. $v\in[0,2\pi[$), each mode is two-fold degenerate. We then considered the half-tractoid (i.e. $v\in[0,\pi[$) and used the following formulae:
@@
Appendix G Nearest Neighbour Spacing Distribution
The Nearest Neighbour Spacing Distribution $P(s)$ counts the number of adjacent eigenvalues which are distant of $s$. Since the mean spacing changes with $x=kR_{0}$, it must be normalized by the Weyl law (see App. F) for a proper comparison with the BGS conjecture [18]. The process is called unfolding [21]. It leads to a mean spacing equal to 1, while the statistical properties of the fluctuations are preserved.
The unfolded eigenvalues are calculated as:
@@ Appendix G Nearest Neighbour Spacing Distribution
Finally the spacings are calculated between each pair of consecutive rescaled eigenvalues:

### v3

@@ Appendix B Fourier transform versus Lomb-Scargle periodogram
The Discrete Fourier Transform (DFT) from discrete times $t_{n}$ to frequency $f$ is defined as
@@
Therefore, $\Delta_{n}=k_{n+1}-k_{n}$ is not constant for the experimental spectra. In the top panel of Fig. 4(b), we plot the Fourier transform of the tractoid experimental spectrum of Fig. 4(a) calculated by the FFT algorithm implemented on Python in the scipy library, which corresponds to the modulus squared of formula (32) assuming a constant sampling $\Delta=(k_{2}-k_{1})/N$. For comparison, in the middle panel of Fig. 4(b), we plot the modulus square of expression (32) using the actual $(k_{n})$, which are not equidistant. Note that the main peak is significantly shifted, revealing an artefact: the raw DFT (top panel) is peaked at a length which is longer than the perimeter and thus not consistent with WGMs, while in the middle panel the length is slightly shorter than the perimeter, in agreement with WGMs.
The Lomb periodogram $P_{LS}(f)$ plotted in the bottom panel of Fig. 4(b) was computed by the lombscargle routine implemented on Python in the scipy library. Its expression is given in [30] and is similar to the formula in the seminal paper by [31]. It is peaked at the same length as the well-sampled DFT in the middle panel of Fig. 4(b), however, the signal-to-noise ratio is much better than for the other two procedures. Therefore, we prefer the Lomb periodogram for analyzing the experimental spectra.
@@ Appendix C Stability computations of periodic geodesics
This Appendix deals with the stability of geodesics on the tractoid surface. First we derive the monodromy matrix for propagation $\mathcal{P}$, then the monodromy matrix for reflection $\mathcal{R}$. Finally in App. C.3 and C.4 we apply these formulas to periodic geodesics of the full and truncated tractoid, respectively.
@@ C.1 Monodromy matrix for propagation
In this paragraph, we derive the monodromy matrix $\mathcal{P}(t)$ for the propagation along a geodesic on the tractoid surface. The usual $(\zeta,\sin(\chi))$ coordinate system along the unperturbed trajectory is convenient, where $\zeta$ is the coordinate in a direction orthogonal to the geodesic tangent vector and $\chi$ is the deviation from the original propagation direction, as shown in Fig. 14. The coordinate $\sin(\chi)$ corresponds to $d\zeta/dt$, ie. the local slope with respect to the unperturbed geodesic. Within this linearized stability analysis, the matrix $\mathcal{P}(t)$ is defined as
$\begin{pmatrix}\zeta(t)\\ \zeta^{\prime}(t)\end{pmatrix}=\mathcal{P}(t)\begin{pmatrix}\zeta(0)\\ \zeta^{\prime}(0)\end{pmatrix}$ (33)
The Jabobi equation [16] describes the evolution of $\zeta$ along a given geodesic as a function of $t$ and reads
@@
Calculating the derivative of Eq. (35) and comparing it with (33) yields
$\mathcal{P}(t)=\begin{pmatrix}\cosh\left(\frac{t}{R_{0}}\right)&R_{0}\sinh\left(\frac{t}{R_{0}}\right)\\ \frac{1}{R_{0}}\sinh\left(\frac{t}{R_{0}}\right)&\cosh\left(\frac{t}{R_{0}}\right)\end{pmatrix}$ (36)
which is the monodromy matrix for propagation along a geodesic on the tractoid. As expected, it has the property $\mathcal{P}(t_{1})\mathcal{P}(t_{2})=\mathcal{P}(t_{1}+t_{2})$ and it tends to the propagation matrix in a planar billiard in the limit of vanishing Gaussian curvature $R_{0}\to\infty$. Note that $\det(\mathcal{P})=1$ and $\operatorname{tr}(\mathcal{P})=2\cosh(\frac{t}{R_{0}})$ which is larger than 2. Therefore, one of its eigenvalues is larger than 1 and the propagation is unstable as it is expected on a surface with a negative curvature.
Figure 14: Schematic for the computation of the propagation monodromy matrix. The black curve is the reference geodesic. It is parametrized by $t$ such that the black dot corresponds to $t=0$. The blue dotted line is the perturbated geodesic. The initial point at $t=0$ is represented by the blue dot and is shifted by a distance $\zeta(0)$. The initial angle of the blue curved with respect to the reference is called $\chi$, which is equal to $\zeta^{\prime}(0)$.
@@ C.2 Monodromy matrix for reflection
In this section, we derive the monodromy matrix for the reflection of a ray on the surface boundary, following Ref. [21]. First, the general case of a given surface and a given boundary is studied. Then this result is applied to the specific case of the tractoid.
@@ C.2.1 The general case
The surface is defined by the vector $\vec{F}(u,v)$ [cf. Eqs. (1-3)] and parameterized by the variables $u$ and $v$. The boundary is a curve belonging to the surface and defined by the vector $\vec{c}\,(t)$. The parametrization is chosen such that $\|\frac{d\vec{c}}{dt}\|=1$. This choice simplifies the definition of the curvature of the curve which is then the modulus of the second derivative: $\kappa=\|\frac{d^{2}\vec{c}}{dt^{2}}\|$.
@@
It remains to calculate the tangent curvature $\kappa_{t}$. To formalize that, we denote $\vec{N}$ the unit vector normal to the surface. If the curvature of the boundary $\frac{d^{2}\vec{c}}{dt^{2}}$ is collinear to $\vec{N}$ then all the curvature is outside the tangent plane and thus the tangent curvature $\kappa_{t}$ is zero. This is the case when the curve $\vec{c}$ is a geodesic. Hence, a reflection on a geodesic is equivalent to a reflection on a plane mirror in the 2D Euclidean case.
In general, the vector normal to the surface $\vec{N}$ and the curvature of the boundary $\vec{c}$ are not collinear. The tangent curvature $\kappa_{t}$ is the modulus of the projection of $\frac{d^{2}\vec{c}}{dt^{2}}$ onto the tangent plane:
@@ C.3 Stability computations for the full tractoid
In this paragraph, the full infinite tractoid is considered for which only the periodic geodesics described in Fig. 5 exist. Using the same notations as in Sec. IV, the total monodromy matrix for the propagation of the $(M,P)$-periodic geodesic can be written
@@
with $\mathcal{M}_{0}(M,P)=\mathcal{P}(M,P)\mathcal{R}(M,P)$ the product of the propagation matrix and the reflection matrix for the $(M,P)$-periodic geodesic. Combining formulas (13) and (14) with the expression of the matrices in Eqs. (15) and (16), we obtain the following expressions for $\mathcal{P}(M,P)$
$\mathcal{P}(M,P)=\begin{pmatrix}\Pi_{1}&R_{0}\Pi_{2}\\ \frac{\Pi_{2}}{R_{0}}&\Pi_{1}\end{pmatrix}$
@@
and for $\mathcal{R}(M,P)$
$\mathcal{R}(M,P)=\begin{pmatrix}-1&0\\ \frac{2}{R_{0}}\frac{M}{P\pi}\left(1+\left(\frac{P\pi}{M}\right)^{2}\right)^{\frac{1}{2}}&-1\end{pmatrix}$ (47)
@@ C.4 Stability computations for the truncated tractoid
However, rays can also be reflected on the top boundary and therefore a new type of periodic geodesic appears. Since the top boundary is also a straight horizontal line, the reflection angles are preserved. It entails that such periodic geodesics are composed of a series of identical patterns, made of circular arcs as illustrated in Fig. 15. Each such periodic geodesic is uniquely identified by a pair of positive relatively prime integers that will also be called $(M,P)$, $M$ being the number of reflections on the lower boundary and $P$ the rotation number. The $(4,3)$ periodic geodesic is sketched in Fig. 15.
@@
As previously, the monodromy matrix of the $(M,P)$ periodic geodesics $\mathcal{M}^{\prime}(M,P)$ can be written
@@
$\mathcal{M}^{\prime}_{0}(M,P)=\mathcal{P}(M,P)\mathcal{R}(M,P)\mathcal{P}(M,P)\tilde{\mathcal{R}}(M,P)$ (52)
where $\mathcal{P}(M,P)$ accounts for the propagation from the upper boundary to the lower one or the opposite, $\mathcal{R}(M,P)$ for the reflection at the boundary $y=1$ and ~$\tilde{\mathcal{R}}(M,P)$ for the reflection at the boundary $y=y_{max}$. Note that $\epsilon=+1$ in $\mathcal{R}(M,P)$, and $\epsilon=-1$ in ~$\tilde{\mathcal{R}}(M,P)$.
After cumbersome but trivial calculations, we obtain that the trace of $M^{\prime}_{0}(M,P)$ is equal to 2. These periodic geodesics are thus also marginally stable as expected. It should be noted that all the periodic geodesics discussed in this article come in families since they are invariant under rotation (that is, a shift of the starting coordinate $v$).
@@ Appendix E Matrices for solving the scalar wave equation
In this Appendix, we provide some elements to solve the scalar wave equation of Sec. V.4 via an eigenvalue problem.
We start by spatially discretizing the problem. We consider $N+2$ equally spaced points $0=\eta_{0}<\eta_{1}<...<\eta_{N}<\eta_{N+1}=\eta_{max}$ and we note $\xi_{i}=\xi(\eta_{i})$ for $i\in[\![0,N+1]\!]$. For a sufficiently small spatial step $h=\frac{\eta_{N+1}-\eta_{0}}{N+1}$, the second order derivative is well approximated by the discretization
for $i\in[\![1,N]\!]$.
The following matrices are then used:
$\mathcal{T}=\frac{1}{h^{2}}\begin{pmatrix}2&-1&0&\dots&0\\ -1&2&-1&&\\ &&\ddots&&\\ &&-1&2&-1\\ 0&\dots&0&-1&2\end{pmatrix}$ (63)
$\mathcal{V}=m^{2}\begin{pmatrix}e^{2\eta_{1}}&&\\ &\ddots&\\ &&e^{2\eta_{N}}\end{pmatrix}$ (64)
@@ Appendix F Weyl law
The average number of eigenmodes $N_{smooth}(x)$ in an energy interval $[0,x]$ with $x:=kR_{0}$ is given by the Weyl law [4],
$N_{Weyl}(x)=\frac{\mathcal{A}}{4\pi}x^{2}-\frac{\mathcal{P}}{4\pi}x-\frac{\mathcal{A}}{12\pi}$ (65)
where $\mathcal{P}$ and $\mathcal{A}$ are the perimeter of the tractoid in units of $R_{0}$, and the area of the tractoid in units of $R_{0}^{2}$, respectively. For the full tractoid (i.e. $v\in[0,2\pi[$), each mode is two-fold degenerate. We consider the half-tractoid (i.e. $v\in[0,\pi[$) and obtain
@@
Appendix G Nearest-Neighbour Spacing Distribution
The Nearest-Neighbour Spacing Distribution $P(s)$ is the probability to find a spacing of two adjacent eigenvalues in an interval $[s,s+ds]$. For a proper comparison with the BGS conjecture [26] which makes statements about universal spectral properties, we need to extract the $k$-dependence of the eigenmode mean density, which is deduced from Weyl’s law (see App. F). This process is referred to as unfolding [32]. It leads to a mean spacing equal to 1, while the statistical properties of the fluctuations are preserved.
The unfolded eigenvalues $\tilde{x}_{i}$ are obtained by replacing the eigenvalues $x_{i}$ by Weyl’s law
@@ Appendix G Nearest-Neighbour Spacing Distribution
Finally, the spacings between adjacent unfolded eigenvalues is given as